CuteAsDucks:

how do you solve x2+6x+10=0 ASAP

CuteAsDucks:

@acehuert

carlosbigbain:

solve the equation x^2 + 6x + 10 = 0 using the quadratic formula. The quadratic formula is: x = (-b ± √(b^2 - 4ac)) / (2a) For our equation, a = 1, b = 6, and c = 10. Plugging in these values, we get: x = (-6 ± √(6^2 - 4*1*10)) / (2*1) Simplifying further: x = (-6 ± √(36 - 40)) / 2 x = (-6 ± √(-4)) / 2 Since we have a negative value inside the square root, the solutions will be complex numbers. Let's simplify it a bit more: x = (-6 ± 2i) / 2 Now we can simplify it even further: x = -3 ± i So the solutions to the equation x^2 + 6x + 10 = 0 are x = -3 + i and x = -3 - i.

carlosbigbain:

Medals are apricated

CuteAsDucks:

Thanks !

acehuert:

Use the quadratic formula\(x=\frac{-\textcolor{#D24040}{b}\pm \sqrt{\textcolor{#D24040}{b}^{2}-4\textcolor{#B14BA5}{a}\textcolor{#3172E0}{c}}}{2\textcolor{#B14BA5}{a}}\)Once in standard form, identify a, b, and c from the original equation and plug them into the quadratic formula.\(x^{2}+6x+10=0\)\(a=\textcolor{#B14BA5}{1}\)\(b=\textcolor{#D24040}{6}\)\(c=\textcolor{#3172E0}{10}\)\(x=\frac{-\textcolor{#D24040}{6}\pm \sqrt{\textcolor{#D24040}{6}^{2}-4\cdot \textcolor{#B14BA5}{1}\cdot \textcolor{#3172E0}{10}}}{2\cdot \textcolor{#B14BA5}{1}}\)Separate the equations To solve for the unknown variable, separate into two equations: one with a plus and the other with a minus.\(x=-3+i\\ x=-3-i\)

toga:

The given equation, x^2 + 6x + 10 = 0, is a quadratic equation. To solve this equation, we can use the quadratic formula, which is: x = [-b ± sqrt(b^2-4ac)] / 2a where a, b, and c are the coefficients of the quadratic equation. Substituting the values in the given equation, we get: a = 1, b = 6, and c = 10 So, we get: x = [-6 ± sqrt(6^2 - 4(1)(10))] / 2(1) Simplifying this equation, we get: x = [-6 ± sqrt(36 - 40)] / 2 x = [-6 ± sqrt(-4)] / 2 x = [-6 ± 2i] / 2 x = -3 ± i Therefore, the solutions of the given quadratic equation are x = -3 + i and x = -3 - i.

CuteAsDucks:

@toga wrote:
The given equation, x^2 + 6x + 10 = 0, is a quadratic equation. To solve this equation, we can use the quadratic formula, which is: x = [-b ± sqrt(b^2-4ac)] / 2a where a, b, and c are the coefficients of the quadratic equation. Substituting the values in the given equation, we get: a = 1, b = 6, and c = 10 So, we get: x = [-6 ± sqrt(6^2 - 4(1)(10))] / 2(1) Simplifying this equation, we get: x = [-6 ± sqrt(36 - 40)] / 2 x = [-6 ± sqrt(-4)] / 2 x = [-6 ± 2i] / 2 x = -3 ± i Therefore, the solutions of the given quadratic equation are x = -3 + i and x = -3 - i.
tyy

toga:

@cuteasducks wrote:
@toga wrote:
The given equation, x^2 + 6x + 10 = 0, is a quadratic equation. To solve this equation, we can use the quadratic formula, which is: x = [-b ± sqrt(b^2-4ac)] / 2a where a, b, and c are the coefficients of the quadratic equation. Substituting the values in the given equation, we get: a = 1, b = 6, and c = 10 So, we get: x = [-6 ± sqrt(6^2 - 4(1)(10))] / 2(1) Simplifying this equation, we get: x = [-6 ± sqrt(36 - 40)] / 2 x = [-6 ± sqrt(-4)] / 2 x = [-6 ± 2i] / 2 x = -3 ± i Therefore, the solutions of the given quadratic equation are x = -3 + i and x = -3 - i.
tyy
np